### v1

# OPDR: Order-Preserving Dimension Reduction for Semantic Embedding of Multimodal Scientific Data

Abstract
One of the most costly operations in scientific data management is searching for the $k$ most similar items (or, $k$-nearest neighbors, KNN) from the database after being provided a new item. The state-of-the-art approach to solving this problem is building an index (e.g., DBSCAN, K-means), such that some tuples can be disqualified at an early stage. However, those existing techniques are inappropriate when the scientific data are multimodal as the measures for different modals are incomparable. The recent advances of multimodal machine learning could potentially overcome the multimodality issue by offering an alternative way for building a semantic index, the so-called embedding vectors mapped from the original multimodal data. However, the resulting (embedding) vectors are usually on the order of hundreds or a thousand of dimensions, which are impractically high for time-sensitive scientific applications.
This work proposes to reduce the dimensionality of the output embedding vectors such that the set of top-$k$ nearest neighbors do not change in the lower-dimensional space, namely Order-Preserving Dimension Reduction (OPDR). In order to develop such an OPDR method, our central hypothesis is that by analyzing the intrinsic relationship among key parameters during the dimension-reduction map, a quantitative function may be constructed to reveal the correlation between the target (lower) dimensionality and other variables. To demonstrate the hypothesis, this paper first defines a formal measure function to quantify the KNN similarity for a specific vector, then extends the measure into an aggregate accuracy of the global metric spaces, and finally derives a closed-form function between the target (lower) dimensionality and other variables. We incorporate the closed-function into popular dimension-reduction methods, various distance metrics, and embedding models. Experiments in multiple scientific data sets show that the proposed OPDR method accurately capture the relationship between the target dimensionality and other variables.
Introduction
Background and Motivation
Many scientific applications involve heterogeneous data resources and cannot directly apply conventional indexing techniques (e.g., DBSCAN, $k$-means) to accelerate query performance. For example, in material sciences (Rangel DaCosta et al. 2021), each material entity comprises multiple pictures along with textual description. Those textural and pictorial (e.g., stored as numerical matrices) data are usually managed by a scientific database where each type of data is managed by a finer-granularity structure, such as tables in relational databases or blobs in NoSQL databases. The conventional wisdom from a data management point of view to efficiently manage those heterogeneous data is to construct a distinct index for each of those structures (e.g., tables, blobs), which does not necessarily reflect the unified, or semantic, measure of the heterogeneous data.
The emergence of multimodal machine learning (e.g., (Li et al. 2022) provides a plausible method to accelerate the queries over heterogeneous scientific data, thanks to the open-source language models that can be directly used for inference; however, the structure of such models cannot be trivially customized and therefore, the potential high dimensionality of the embedding vectors, usually on the order of hundreds or thousands, leads to a problem known as the “curse of dimensionality.” This issue precludes multimodal machine learning from broad deployment due to the high-performance (i.e., time-sensitive) nature of scientific applications. For example, many language models (BERT  (Devlin et al. 2018), ViT (Dosovitskiy et al. 2020), etc.) generate the embedding vectors of 512 or more dimensions; however, scientific applications (Rangel DaCosta et al. 2021; Wu et al. 2020) require to analyze billions of data (i.e., vectors) on the order of milliseconds. The curse of dimensionality exacerbates the computational complexity, not to mention the additional challenges for model accuracy and generalization.
There is therefore an urgent need to develop new methods to fill the gap between the high dimensionality of embedding vectors generated by pretrained language models and the desired lower dimensionality, and this work focuses on one of the most common analytic tasks in scientific applications—$k$-nearest neighbors (KNN) queries. The KNN query is almost ubiquitous in scientific applications due to the nature of scientific discoveries: In many disciplines, especially natural sciences (Lu et al. 2023; Tshitoyan et al. 2019; Gupta et al. 2022), matching a newly found/constructed item to another well-studied entity in the database is one of the most effective methods to study the new item. If the new method could significantly reduce the performance overhead for performing the KNN queries, many scientific applications would benefit from the new method.
Proposed Work
In order to reduce the dimension of high-dimensional embedding vectors in a meaningful sense, this paper first proposes a new notion, namely Order-Preserving Measure (OPM), for the set of nearest neighbors that is preserved during a dimension-reduction map between two metric spaces. In essence, the OPM is formalized as a mathematical function that is well defined over the power-set $\sigma$-algebra over the lower-dimensional metric space. This allows us to assign a numerical (real and non-negative) value to an arbitrary subspace of the low-dimensional space and consequently, apply aggregation to quantify the overall accuracy of the dimension-reduction map in terms of preserving the set of $k$-nearest neighbors of each embedding vector.
Built upon the aforementioned OPM, this paper then constructs a global metric to quantify the closeness of two metric spaces $X$ and $Y$ in terms of, again, the preservation of the set of $k$-nearest neighbors from $X$ to $Y$. The objective of constructing such a global metric is to derive a closed-form function among important environmental parameters, such as the dimensionality of metric spaces, the number of data points, and the expected closeness, or accuracy, between the metric spaces in terms of $k$-nearest neighbors.
The conjectured closed-form function is constructed through multiple observations and qualitative analysis. The expected dimension of the lower-dimensional space is positively correlated to both the space’s dimensionality and cardinality (i.e., the number of embedding vectors), but with different weights. Specifically, we assign significantly higher weight (exponentiation) on the closeness metric (also known as accuracy, which will be formally defined) than that on the cardinality.
In addition to theoretical results about the closed-form function, this paper also demonstrates how to incorporate the order-preserving measure and the closed-form function into the workflow of real-world scientific applications. We will detail the implementation of both the OPM and the closed-form function, and demonstrate that how they work together with other subsystems, such as dimension-reduction methods (PCA, MDS), neural network models (CLIP, ViT, Bert), and distance metrics (Euclidean, cosine, Manhattan). We then evaluate the augmented dimension-reduction system with four multimodal scientific datasets of representative characteristics; our results show that the function accurately captures the intrinsic relationship among the key factors of a dimension-reduction map and therefore, contributes new insights to leveraging advanced AI techniques for complex scientific applications.
Contributions
In summary, this work makes the following scientific and technical contributions.
  • We present a new notion for dimension reduction, namely Order-Preserving Dimension Reduction (OPDR), which leads to a more flexible metric that suffices to capture the invariant of $k$-nearest neighbors.
  • We construct a closed-form function that captures the intrinsic relationship among the target dimensionality and other factors in the context of OPDR.
  • We implement the proposed OPDR method and demonstrate the effectiveness of OPDR by carrying out an extensive evaluation on real-world multimodal scientific data.
Preliminaries and Related Work
Multimodal Scientific Data
Multimodal scientific data has drawn a lot of research interests. For instance, the MELINDA dataset (Wu et al. 2020) focuses on classifying biomedical experiment methods by utilizing multiple data types. Similarly, a study on explainable AI (Jin, Li, and Hamarneh 2022) evaluates algorithms within the context of multimodal medical imaging tasks, emphasizing the importance of decision-making based on raw scientific data.
Researchers have also trained new large language models specifically designed for scientific data. For example, GIT-Mol (Liu et al. 2024) represents a notable effort, developing a multimodal large language model for molecular science that integrates graph, image, and text data. Additionally, MatSciBERT (Gupta et al. 2022) focuses on creating domain-specific language models tailored to materials science. While these models, along with others such as the unsupervised word embeddings for materials science literature (Tshitoyan et al. 2019), demonstrate the potential of specialized models in capturing domain-specific knowledge, they are often limited in scope, typically focusing solely on text-based data.
Multi-modal data is often highly dimensional, making it crucial to find a trade-off between maintain essential information and reducing the data’s complexity. Therefore, some scientists have proposed effective methods to address this challenge. For instance, the study on Multiscale Feature Extraction and Fusion of Image and Text in VQA (Lu et al. 2023) introduces advanced techniques to fuse multimodal data, allowing for a more integrated analysis of different data types. The research on Semi-Supervised Multi-Modal Learning with Balanced Spectral Decomposition (Hu et al. 2020) focuses on optimizing information preservation through spectral decomposition, which enhances the ability to capture correlations within the data. Additionally, Unsupervised word embeddings in materials science literature (Tshitoyan et al. 2019) demonstrates the transformation of textual information into vectors, providing a compact and efficient representation of large-scale text data.
Contrastive Language-Image Pre-Training
With the introduction of “Attention is All You Need” (Vaswani et al. 2017), the development of large language models has surged, leading to significant advancements in the field. Following this breakthrough, models like BERT (Devlin et al. 2018) and ViT (Dosovitskiy et al. 2020) have focused on text analysis and image recognition respectively. Particularly noteworthy is the CLIP model (Radford et al. 2021), which integrates both text and image data, marking a significant step forward in multimodal data processing. All kinds of studies that improve CLIP model has been propsed. For instance, TiMix (Jiang et al. 2024), address issues with noisy web-harvested text-image pairs by using mixed data samples, while CLIP-Event (Li et al. 2022) and SoftCLIP (Gao et al. 2024) further enhance the alignment between text and images. In (Zhang et al. 2021), authors proposed to convert images into binary codes (hash codes) that preserve image similarity, making it easier to compare images within large-scale datasets.
Some prior works proposed methods to perform dimension reduction over the embedding vectors. For example, in (Chen et al. 2022), authors proposed a new approach to reducing dimensionality following contrastive learning. Similarly, in (Huang et al. 2020), authors focused on relative positioning in data before and after LLM manipulation. However, none of the above works touched on the preservation of the set of $k$-nearest neighbors.
Dimension Reduction
After Multidimensional Scaling (MDS) (Torgerson 1952) was published, numerous methods were proposed that preserved pairwise distances in lower-dimensional spaces. For example, Colored Maximum Variance Unfolding (Song et al. 2007) preserved local distances while maximizing variance, and Neighborhood Preserving Embedding (He et al. 2005) retained local relative positions. Further innovations include Tensor Embedding Methods (Dai and Yeung 2006), which tackled the curse of dimensionality, and MultiMAP (Jain et al. 2023), which integrated multimodal data. Recent approaches like Similarity Order Preserving Discriminant Analysis (Hu, Feng, and Chen 2021) and Ordinal Data Clustering Algorithm with Automated Distance Learning (Zhang and Cheung 2020) specifically aimed to preserve data order post-reduction, highlighting the evolving focus on maintaining data integrity across various dimensions. However, none of the above works considered preserving the set of $k$-nearest neighbors during the dimension reduction.
As large language models (LLMs) become popular, scientists have also proposed methods that integrate LLMs with dimension-reduction techniques to enhance the analysis of complex data. For example, Transformer-based Dimensionality Reduction (Ran, Gao, and Fang 2022) introduced a method that decomposes autoencoders into modular components, leveraging the power of transformers to manage high-dimensional data efficiently. Similarly, in (George and Sumathy 2023), authors combined BERT with dimensionality reduction to enhance topic modeling by clustering textual data more effectively. However, the above works assumed that the users of LLMs were able to modify the model structure, which is unrealistic in many scientific applications.
Order-Preserving Dimension Reduction
This section first presents a new notion, i.e., Order-Preserving Measure (OPM), for the set of $k$-nearest neighbors that are preserved during a map between two metric spaces. Then, a closed-form function is constructed to quantify the relationship among the space dimensionality, the number of data points, and the OPM. Finally, a detailed implementation is presented about how to incorporate the new measure and function into scientific applications.
Order-Preserving Measure
Intuitively speaking, an order-preserving measure (OPM) provides a metric of the number of nearest neighbors that do not change between two spaces. For example, if the $2$-closest points (we assume that the spaces are metric spaces such that pair-wise distances are well defined) of $every$ point in a metric space $(X,\delta_{X})$ are still the same $2$-closest points (of each point) in a metric space $(Y,\delta_{Y})$, then we say the map $f:X\to Y$ is order-preserving of 2, or $OP_{2}$. We will provide a more formal definition of $OP_{z}$, $z\in\mathbb{Z}^{+}\cup\{0\}$ later; before that, we need to point out a common misunderstanding of this notion regarding inclusiveness, as follows.
[…]
Mathematically speaking, a measure is a function, say $\mu$, which maps a subset $E_{i}$ of set $X$ in a well-defined $\sigma$-algebra to a value in the extended real line $[0,\infty]$, such that (i) $\mu(\emptyset)$ = 0 and (ii) $\mu\left(\bigcup_{i=1}^{n}E_{i}\right)=\sum_{i+1}^{n}\mu(E_{i})$. Because we assume there exists a well-defined $\sigma$-algebra over (the set of) embedding vectors, say $X$, $E_{i}$’s will be understood as disjoint subsets of $X$ unless otherwise stated. In fact, we will explicitly construct the $\sigma$-algebra on the target space, $Y$. As a side note, a measure function $\mu$ can be also thought of as a homomorphism between the two monoids (i.e., groups without inverses) $(\mathcal{P}(X),\cup)$ and (+$(\mathbb{R}^{+}\cup\{0\},+)$ with kernel $Ker_{\mu}=\{\emptyset\}\subseteq\mathcal{P}(X)$, where $\mathcal{P}(X)$ denotes the power set of $X$. However, our analysis in the following will not require any group-theoretical results.
The construction of the $\sigma$-algebra of embedding vectors is as follows. Let $k$ denote the target number of preserved $k$-nearest neighbors. Let $Y$ denote the target (i.e., low-dimensional) metric space (we omit the metric functions here) of the dimension-reduction map. In practice, the cardinality of $Y$ is finite (i.e., there is an upper limit of number of embedding vectors in scientific applications), which is of course countable. The $\sigma$-algebra $\mathcal{M}_{Y}$ of $Y$ is simply the power set of all the mapped vectors in $Y$, i.e., $\mathcal{M}_{Y}=\mathcal{P}(Y)$, which is obviously a $\sigma$-algebra on $Y$ because $Y$ is countable and $\mathcal{P}(Y)$ is closed for any countable number of set unions.
After constructing its $\sigma$-algebra, we are ready to define the measure on $\mathcal{M}_{Y}$. Let $E^{Y}_{k,i}$ denote the set of the $k$-nearest neighbors of $y_{i}$, where $y_{i}\in Y$. The same notation can be defined for space $X$, which denotes the original metric space where the original vectors of $y_{i}$ reside, i.e., $x_{i}=f^{-1}(y_{i})$, where $f$ denotes the dimension-reduction function. Let $\mu$ be a function defined as follows,
$\mu(F_{j})=\frac{\left|F_{j}\cap E^{Y}_{k,i}\cap E^{X}_{k,i}\right|}{k},$ (1)
where $F_{j}\in\mathcal{M}_{Y}$. It should be clear that $\mu\in[0,1]\subset[0,+\infty]$ because $\left|E^{Y}_{k,i}\right|\in[0,k]$, where ||$\cdot$|| ⋅ | denotes the cardinality of a set.
We argue that the function defined in Eq. (1) is a measure on $\mathcal{M}_{Y}$. This means that we need to demonstrate two properties of $\mu$ (on $\mathcal{M}_{Y}$): (i) $\mu(\emptyset)=0$ and (ii) $\mu(\bigcup_{i=1}^{\infty}F_{j})=\sum_{i=1}^{\infty}(\mu(F_{j}))$, $\forall F_{j}\subseteq Y$ and $F_{j}$’s are disjoint. Property (i) is trivially satisfied because when $F_{j}=\emptyset$, we have $\mu(F_{j})=\frac{\left|\emptyset\right|}{k}=\frac{0}{k}=0$. To prove property (ii), it suffices to show that if $F_{j}=F_{1}\cup F_{2}$ and $F_{1}\cap F_{2}=\emptyset$, then $\mu(F_{1}\cup F_{2})=\mu(F_{1})+\mu(F_{2})$; this is because $\mathcal{M}_{Y}$ is finite and the above binary relationship can be extended finite times. To simply the notation, let $E=E^{Y}_{k,i}\cap E^{X}_{k,i}$. It follows that $\mu(F_{j})$ = $\frac{\left|(F_{1}\cup F_{2})\cap E\right|}{k}$ = $\frac{\left|(F_{1}\cap E)\cup(F_{2}\cap E)\right|}{k}$. Now, because $F_{1}\cap F_{2}=\emptyset$, we know $(F_{1}\cap E)\cap(F_{2}\cap E)=\emptyset$. This implies that there is no common element between $(F_{1}\cap E)$ and $(F_{2}\cap E)$, which means the sets are separate components and the cardinality of the union set is simply the addition of cardinality of each component: $\left|(F_{1}\cap E)\cup(F_{2}\cap E)\right|=\left|F_{1}\cap E\right|+\left|F_{2}\cap E\right|$. It follows that $\mu(F_{1}\cup F_{2})=\frac{\left|F_{1}\cap E\right|+\left|F_{2}\cap E\right|}{k}=\frac{\left|F_{1}\cap E\right|}{k}+\frac{\left|F_{2}\cap E\right|}{k}=\mu(F_{1})+\mu(F_{2})$, as desired.
The discussion about preserving the set of $k$-nearest neighbors between metric spaces around Eq. (1) is, in our humble opinion, a tip of the iceberg regarding the neighborhood-preserving dimension-reduction maps. The reasoning is that the $k$-nearest neighbors can be thought of the discrete representation of the neighborhood of the referred point, or vector. If we extend the idea to a continuous counterpart, i.e., given a point $p$ and its neighborhood or an open set around it, say $\epsilon(p)$, the question becomes under what conditions the open set of the mapped vectors, namely $f(p)$, corresponds to the open set of $p$ in a dimension-reduction function $f$. In topology and analysis, such a function carrying the open sets in the forward direction is called an open map, which is unfortunately not the same (in fact, the inverse) condition under which the function is continuous. We leave this as an open question to the community.
Closed-Form Function
The previous section defines an additive measure $\mu$ on the $\sigma$-algebra of the projected space $Y$; This section investigates the relationship between $\mu$ and other parameters, such as the space dimensionality and the number of data points, or space cardinality. The idea is that we will construct a function for the above variables that we conjecture are critical during the dimension reduction, largely based on our experiment observations and mathematical intuition. We will later in the evaluation section verify the constructed function as our hypothesis.
We assume that we have an existing dimension-reduction method on hand and we are interested in quantifying the target dimensionality of the lower-dimensional space in which the set of $k$-nearest neighbors is an invariant, as defined in Eq. (1). It turns out that before attempting to solve the problem, we need to have a more “global” metric than $\mu(\cdot)$, which works as a local measure that is concerned with only the $k$-nearest neighbors of a single point.
We define the global metric for quantifying the overall closeness (or, similarity, accuracy) between two metric spaces regarding their $k$-nearest neighbors as follows. We first aggregate the measure for each point, then normalize the aggregate measure (AM) into $[0,1]$, and finally calculate the arithmetic mean of all normalized AMs (NAMs). We call the above average of NAMs as the accuracy of $Y$ with respect to $X$ on $k$-nearest neighbors. Formally, we define the accuracy $A$ as
$\text{A}^{X}_{k}(Y)=\frac{1}{m}\cdot\sum_{i=1}^{m}\frac{\mu_{i}\left(Y\setminus\{y_{i}\}\right)}{k},$ (2)
where $m=\left|Y\right|=\left|X\right|$, $y_{j}\in Y$, and $u_{i}(\cdot)$ is the measure on $Y$ over $y_{i}$ as defined in Eq. (1). It is not hard to see that the accuracy $A$ defined above falls within the range $[0,1]$, because $\mu_{i}\left(Y\setminus\{y_{i}\}\right)$ is bounded by $k$ (inclusively).
The parameters in our hypothetical function include the accuracy $A_{k}$ between two metric spaces $X$ and $Y$, the number of data points $m=\left|X\right|=\left|Y\right|$, and the dimensionality of the spaces. In practice, the dimension of the domain space $X$ is determined by the neural network model; therefore, our function would assume that $X$’s dimension is a constant and only involve the dimension of $Y$, denoted by $dim(Y)$. Thus, we expect to construct a function $g$ as follows,
$dim(Y)=g\left(A_{k},m\right),$
such that the result $dim(Y)$ can be set as a parameter in a dimension-reduction function $f$. It follows that the real-world users can simply compose the functions $g$ and $f$, i.e., $f\circ g$, to ensure that the set of $k$-nearest neighbors is an invariant between two spaces $X$ and $Y$.
The construction of $g$ is as follows. We conjecture the form of $g$ based on the following observations.
  • Firstly, $dim(Y)$ is positively influenced by both the accuracy $A_{k}$ and the cardinality $m$. That is, if the user expects to maintain a higher accuracy of the $k$-nearest neighbors in a lower-dimensional space $Y$, then the target dimension $dim(Y)$ should be also set higher. In the extreme case, if $Y=X$, then $A_{k}=1.0$. Similarly, if there are a large quantity of data points, then the intuition is that we may need to keep a large number of dimensions in the target low-dimensional space $Y$.
  • Secondly, the impact of the accuracy $A_{k}$ should be higher than that of the cardinality $m$. This can be understood with the intuition that $dim(Y)$ is quite sensitive to $A_{k}$ because $A_{k}$ is a global metric of the entire space $Y$. On the other hand, a small change of the number of data points may or may not significantly change the distribution of the space.
Based on the above discussion, we conjecture that the function $g$ takes the following form
$g(a,b)=\mathcal{O}(b\cdot 2^{a}),$ (3)
where the exponential factor $2^{a}$ underscores the significantly higher impact of $a$. In other words, the closed-form function is in the form of
$dim(Y)=\mathcal{O}(m\cdot 2^{A_{k}}).$
If we simplify the asymptotic notation with constant coefficients, we can also rewrite the function as
$A_{k}=c_{0}\cdot\log\frac{dim(Y)}{m}+c_{1},$ (4)
where $c_{0}$ and $c_{1}$ are two constant values that can be estimated by various regression models. When $A_{k}=1$, we say the low-dimensional space $Y$ is $OP_{k}$ to the original space $X$, which formalizes our previous notion of $OP_{K}$.
In the following sections, we will verify the effectiveness of the function defined in Eq. (4); but before that, we will describe how we implement the measure $\mu(\cdot)$ and the dimension-reduction function $f$ guided by Eq. (4).
Incorporation into Scientific Data Management
We will first describe how the embedding vectors are generated from multimodal scientific data and then present how we implement the hypothetical function and incorporate it into popular dimension-reduction methods.
To convert scientific data into embedding vectors using CLIP (Contrastive Language-Image Pretraining), Vision Transformer (ViT), and BERT (Bidirectional Encoder Representations from Transformers), which are all transformer-based models, we begin by collecting and preparing the scientific data, which includes both textual data from HDF5 files and images in formats such as TIFF, PNG, or JPEG. The textual data is extracted by reading the Hierarchical Data Format (HDF5) files, where the dataset names and contents are concatenated into a single string representation. For instance, the HDF5 file used in our study contains a dataset with dimensions of (139, 139, 92) along with supporting metadata, providing a structured representation of the scientific data.
For the image data, each image file is read and converted to RGB format. The images, including results obtained at four different resolutions: HAADF, BF, ADF, and ABF, are found to have consistent dimensions of (369, 369) pixels. Additionally, a summary image with dimensions of (941, 905) pixels was included. Then, we generate embeddings for both textual and image data using the CLIP, ViT, and BERT models. For textual data, we utlized BERT model to generate embbedings, which processes the concatenated textual information. Image embeddings are produced by the ViT model, which processes the image files. Then, we utilized CLIP model to analyze both textual and image data.
The generated embeddings maintain their default dimensionality, which is specific to each model: BERT produces embeddings of 768 dimensions, while ViT also outputs 768-dimensional embeddings. For CLIP, the textual data is embedded via CLIP’s text encoder, resulting in 512-dimensional embeddings. Similarly, the image data is processed through CLIP’s image encoder, also yielding 512-dimensional embeddings. To create a comprehensive representation of the multimodal scientific data, we combined these embeddings by concatenating the 512-dimensional text and image vectors into a single, unified embedding vector of 1024 dimensions. All generated embeddings, derived from both textual data and image data at various resolutions, are collected and stored for further analysis.
Following the embedding vector extraction, we focus on the dimensionality reduction phase. We evaluated several mainstream techniques including PCA, and MDS, among which PCA consistently outperformed others in terms of maintaining the integrity of location information in our datasets. Our investigations revealed a notable correlation between the effectiveness of PCA in preserving the spatial relationships and the ratio $\frac{n}{m}$, where $n=dim(Y)$ and $m=\left|Y\right|$, of the target dimension to the number of samples used in the PCA process. We adopted various regression models to elucidate the relationship between the accuracy of preserved relative location information and the ratio $\frac{n}{m}$. These models facilitate the prediction of an optimal embedding vector dimension required to achieve a predetermined accuracy level, given a known number of samples, $m$.
Evaluation
This section evaluates the effectiveness of the proposed method for KNN-preserved dimension reduction on multiple scientific data sets. To ensure the robustness of our method across various scenarios, we evaluate the method with three Transformer-based models for extracting embedding vectors, two popular dimension-reduction techniques, and three distance metrics. All results suggest that the proposed method is highly effective.
Experimental Setup
Data Sets
The dataset used for evaluation were sourced from the Material Project  (The Materials Project Accessed 2024) and categorized into four distinct datasets: experimental Observable, stable, metal, magnetic. To be more specific, experimental observable dataset includes 33,990 data points, stable dataset consists of 48,884 data points, metal dataset has 72,252 data points, and magnetic dataset consists of 81,723 data points.
Platform
For our experiments, we chose CloudLab (Duplyakin et al. 2019) as the testing platform. In this experiment, the specific machine specifications we selected are as follows: All of our experiments are carried out on a CloudLab machine with the following specifications: Node Type: c6420. The machine is equipped with two sixteen-core Intel Xeon Gold 6142 CPUs at 2.6 GHz, 384 GB ECC DDR4-2666 Memory, and two Seagate 1TB 7200 RPM 6G SATA hard drives. The network interface card (NIC) is a dual-port Intel X710 10GbE. We have installed the following libraries in conda environment: transformer, sklearn, torch, pandas, numpy, os, pickle, h5py, shutill, matlotlib, and pyprismatic.
OPDR on Various Data Sets
We first extracted embedding vectors from four datasets with distinct characteristics (Observable, Unstable, Metal, Non-magnetic) using the CLIP model. Subsequently, we computed the distances between vectors using the L2-norm and performed dimensionality reduction using PCA. To simplify notation, we let $n=dim(Y)$, denote the dimension of the lower-dimensional space. For these four groups of datasets, we further divided them into eight subsets each, with sample sizes $m$ where $m\in\{10,20,30,40,50,60,70,80\}$, to test whether the distribution patterns of data points were consistent across different sample sizes. Observations from figure 4 to figure 4 indicate a strong positive correlation between accuracy and the ratio of $n$ to $m$. As $n$ approaches $m$, accuracy initially increases rapidly and then slows down, converging to a stable value. Although the four different datasets exhibit highly similar data distributions, minor differences may be attributed to the inherent distribution characteristics and randomness of the samples.
Figure 1: Observable
Figure 2: Unstable
Figure 3: Metal
Figure 4: Non-magnetic
Through comparative analysis of the four data sets discussed above, we observe that our proposed OPDR algorithm performs consistently across multiple data sets and exhibits the anticipated data patterns, validating our initial hypotheses. In the remainder of this section, we will focus on the Observable Material dataset, exploring the effects of using different Transformer-based models, dimensionality reduction techniques, and various distance metrics.
Influence of Embedding Models
We continued our experiments on the ‘Observable’ dataset, merely changing the model used for extracting embedding vectors. We employed three distinct transformer-based models: BERT (Bidirectional Encoder Representations from Transformers), Vision Transformer (ViT), and CLIP (Contrastive Language-Image Pre-training) for processing and analyzing scientific data. Specifically, the BERT model was used to analyze text data extracted from HDF5 files, ViT was utilized for processing image data in TIFF and PNG formats, and CLIP was applied to analyze comprehensive multimodal datasets. This enabled us to achieve more holistic and unified representations of scientific data.
Through multiple experiments with these three models, we obtained data fit lines for each model to represent the relationship between accuracy and the ratio $n/m$. All three fit lines indicated that as $n$ approaches $m$, our accuracy first increases rapidly and then converges slowly. The results from the three different models were consistent with our initial hypotheses. Furthermore, we observed that the obtained fit lines were essentially overlapping, indicating that the choice of model does not alter the data structure. Utilizing these diverse transformer-based models, we were still able to ensure the accuracy and consistency of data relationships. Using different models does not change the data format; thus, in practical applications, it is essential to consider the properties of the data and its intended use to select the most appropriate model for experimentation.
Figure 5: Data Fitting for Different Embedding Models
Figure 5 shows that using different models does not alter much the accuracy. This implies that the closed-form function is applicable to a variety of neural network models. Therefore, in practical applications, one should consider the nature of the data and its intended use to select the most appropriate model for experimentation. The original data from which the regression line is calculated can be found in Figure 8 in the Appendix.
Influence of Dimension-Reduction Methods
The scientific data analyzed in this study exhibits extremely high dimensionality, often reaching up to hundreds of thousands of dimensions. Despite the application of advanced embedding techniques, the resulting vectors still contain several hundred to several thousand dimensions, posing considerable challenges for effective data analysis and interpretation. To tackle these issues, we employed two principal dimensionality reduction methods—Multidimensional Scaling (MDS) and Principal Component Analysis (PCA)—both renowned for their ability to efficiently reduce data complexity while maintaining the integrity of the relationships among data points. Given that the majority of contemporary dimensionality reduction techniques build on the foundational theories of PCA and MDS, our study specifically focused on these methods to evaluate their practical efficacy. By implementing MDS and PCA, we aimed to simplify the dataset in a way that preserves the relative positioning of data points, thereby facilitating more manageable and interpretable analysis.
By comparing the performance of PCA and MDS on the same dataset in Figure 6, we observe that PCA is more sensitive to changes in $n/m$ and can reach higher accuracy more quickly. Moreover, PCA can achieve a maximum accuracy of 100%, surpassing the peak accuracy attainable by MDS. This demonstrates PCA’s superior performance in preserving the relative positions of data points. Additionally, employing different dimensionality reduction techniques does not alter the overall data pattern; the relationship between accuracy and $n/m$ still aligns with our hypothesis.
Figure 6: Data Fitting for Different Dimension-Reduction Methods
From Figure 6, although the overall data pattern remains constant and both PCA and MDS align with our prior hypotheses, the choice of dimensionality reduction technique is found to have a noticeable effect on the fitting outcomes. In practical applications, it is crucial to weigh the specific contexts and experimental data at hand. This assessment will inform the selection of the most appropriate dimensionality reduction method to achieve the desired level of accuracy. The original data from which the regression line is calculated can be found in figure 9 in the Appendix.
Influence of Distance Metrics
Figure 7: Data Fitting for Different Distance Metrics
In this section, we explore the impact of using different distance metrics—namely L1 norm, L2 norm, and cosine distance—on the Observable dataset and CLIP model. The geometric meanings of these metrics are distinct:
  • L1 norm (Manhattan distance): Measures the sum of the absolute differences of their coordinates. It represents a very intuitive distance metric, similar to the shortest path between two points in a grid-like structure.
  • L2 norm (Euclidean distance): Calculates the straight-line distance between two points. It is the most commonly used distance metric, applicable in a wide range of natural and engineering problems.
  • Cosine distance: Measures the cosine of the angle between two vectors, reflecting the difference in direction rather than magnitude, often used to gauge similarity in text or high-dimensional data.
Experimental observations shown in Figure 7 revealed that as $n/m$ increases, the behaviors of L2 norm and cosine distance are similar, whereas L1 norm performs slightly worse than the other two. Despite the change in distance metrics not altering the overall data pattern—they still show the trend proposed in Eq (4)—the choice of distance metric remains crucial for specific scenarios.In practical use, experimenters should consider the application scenarios of their data to decide which distance calculation method to use. The original data from which the regression line is calculated can be found in figure 10 in the Appendix.
Conclusion and Future Work
This paper solves an open problem in multimodal scientific data management regarding the high dimensionality of embedding vectors generated by large language models. Theoretical results include new notions (both point-wise and space-level) to measure the closeness of two metric spaces in terms of their $k$-nearest neighbors and a closed-function to reveal the relationship between space dimensionality, space cardinality, and space similarity (or, accuracy of the dimension-reduction map). The closed-form function is implemented and incorporated into other subsystems, such as dimension-reduction methods (PCA, MDS), neural network models (CLIP, ViT, Bert), and distance metrics (Euclidean, cosine, Manhattan), which collectively enable a new dimension-reduction approach for efficiently manage and analyze multimodal scientific data.
Our future work will dive deeper into the theoretical foundation of the closed-form function among metric spaces’ dimensionality, cardinality, and accuracy of the dimension-reduction map; the fact that the proposed order-preserving measure is a well-defined measure may shed more insight than we originally thought. We also plan to investigate the extensibility of the proposed method for production vector database systems, such as PostgreSQL and pgvector.
Appendix A Additional Experimental Results
According to Figure 8, the data distributions obtained from the three Transformer-based models are highly similar. In contrast, the distribution from the BERT model is more dispersed, while the distributions from CLIP and ViT models are more concentrated. The results indicate that the CLIP model performs the most stably, with the overall patterns of the three models being consistent, aligning with our initial hypotheses. The fitting lines for the three groups of data points have been discussed previously in Figure 5.
Figure 8: Original Data for Neural Network Models
Figure 9 clearly shows the accuracy distributions for the ratio $n/m$ using two different dimensionality reduction techniques, PCA and MDS. Both PCA and MDS exhibit highly concentrated data points. Comparatively, PCA demonstrates higher accuracy as $n$ approaches $m$, indicating that PCA has stronger performance capabilities. Overall, the data patterns of both methods are similar, aligning with our initial hypotheses. The fitting lines for these two dimensionality reduction methods have already been discussed in figure 6.
Figure 9: Original Data for PCA and MDS
Figure 10 clearly illustrates the accuracy distributions for the ratio $n/m$ using different distance metrics. Data points generated by the L2-norm and cosine distance highly overlap, indicating similar performance of these metrics. In contrast, the L1-norm shows lower sensitivity and fails to reach the precision of other methods. Overall, the three datasets exhibit similar patterns: accuracy rapidly increases as $n$ approaches $m$ and then slowly converges to a peak. This observation aligns with our initial hypothesis. The fitting lines for the three sets of data points obtained using different distance metrics have been discussed in figure 7.
Figure 10: Original Data for Distance Metrics

### v2

@@
-OPDR: Order-Preserving Dimension Reduction for Semantic Embedding of Multimodal Scientific Data
+Order-Preserving Dimension Reduction for Multimodal Semantic Embedding
 Abstract
-One of the most costly operations in scientific data management is searching for the $k$ most similar items (or, $k$-nearest neighbors, KNN) from the database after being provided a new item.
-The state-of-the-art approach to solving this problem is building an index (e.g., DBSCAN, K-means), such that some tuples can be disqualified at an early stage.
-However, those existing techniques are inappropriate when the scientific data are multimodal as the measures for different modals are incomparable.
-The recent advances of multimodal machine learning could potentially overcome the multimodality issue by offering an alternative way for building a semantic index, the so-called embedding vectors mapped from the original multimodal data.
-However, the resulting (embedding) vectors are usually on the order of hundreds or a thousand of dimensions, which are impractically high for time-sensitive scientific applications.
-This work proposes to reduce the dimensionality of the output embedding vectors such that the set of top-$k$ nearest neighbors do not change in the lower-dimensional space, namely Order-Preserving Dimension Reduction (OPDR).
-In order to develop such an OPDR method, our central hypothesis is that by analyzing the intrinsic relationship among key parameters during the dimension-reduction map, a quantitative function may be constructed to reveal the correlation between the target (lower) dimensionality and other variables.
-To demonstrate the hypothesis, this paper first defines a formal measure function to quantify the KNN similarity for a specific vector, then extends the measure into an aggregate accuracy of the global metric spaces, and finally derives a closed-form function between the target (lower) dimensionality and other variables.
-We incorporate the closed-function into popular dimension-reduction methods, various distance metrics, and embedding models.
-Experiments in multiple scientific data sets show that the proposed OPDR method accurately capture the relationship between the target dimensionality and other variables.
+Searching for the $k$-nearest neighbors (KNN) in multimodal data retrieval is computationally expensive, particularly due to the inherent difficulty in comparing similarity measures across different modalities.
+Recent advances in multimodal machine learning address this issue by mapping data into a shared embedding space; however, the high dimensionality of these embeddings (hundreds to thousands of dimensions) presents a challenge for time-sensitive vision applications.
+This work proposes Order-Preserving Dimension Reduction (OPDR), aiming to reduce the dimensionality of embeddings while preserving the ranking of KNN in the lower-dimensional space.
+One notable component of OPDR is a new measure function to quantify KNN quality as a global metric, based on which we derive a closed-form map between target dimensionality and key contextual parameters.
+We have integrated OPDR with multiple state-of-the-art dimension-reduction techniques, distance functions, and embedding models; experiments on a variety of multimodal datasets demonstrate that OPDR effectively retains recall high accuracy while significantly reducing computational costs.
 Introduction
 Background and Motivation
-Many scientific applications involve heterogeneous data resources and cannot directly apply conventional indexing techniques (e.g., DBSCAN, $k$-means) to accelerate query performance.
-For example, in material sciences (Rangel DaCosta et al.
-2021), each material entity comprises multiple pictures along with textual description.
-Those textural and pictorial (e.g., stored as numerical matrices) data are usually managed by a scientific database where each type of data is managed by a finer-granularity structure, such as tables in relational databases or blobs in NoSQL databases.
-The conventional wisdom from a data management point of view to efficiently manage those heterogeneous data is to construct a distinct index for each of those structures (e.g., tables, blobs), which does not necessarily reflect the unified, or semantic, measure of the heterogeneous data.
-The emergence of multimodal machine learning (e.g., (Li et al.
-2022) provides a plausible method to accelerate the queries over heterogeneous scientific data, thanks to the open-source language models that can be directly used for inference; however, the structure of such models cannot be trivially customized and therefore, the potential high dimensionality of the embedding vectors, usually on the order of hundreds or thousands, leads to a problem known as the “curse of dimensionality.” This issue precludes multimodal machine learning from broad deployment due to the high-performance (i.e., time-sensitive) nature of scientific applications.
-For example, many language models (BERT  (Devlin et al.
-2018), ViT (Dosovitskiy et al.
-2020), etc.) generate the embedding vectors of 512 or more dimensions; however, scientific applications (Rangel DaCosta et al.
-2021; Wu et al.
-2020) require to analyze billions of data (i.e., vectors) on the order of milliseconds.
-The curse of dimensionality exacerbates the computational complexity, not to mention the additional challenges for model accuracy and generalization.
-There is therefore an urgent need to develop new methods to fill the gap between the high dimensionality of embedding vectors generated by pretrained language models and the desired lower dimensionality, and this work focuses on one of the most common analytic tasks in scientific applications—$k$-nearest neighbors (KNN) queries.
-The KNN query is almost ubiquitous in scientific applications due to the nature of scientific discoveries: In many disciplines, especially natural sciences (Lu et al.
-2023; Tshitoyan et al.
-2019; Gupta et al.
-2022), matching a newly found/constructed item to another well-studied entity in the database is one of the most effective methods to study the new item.
-If the new method could significantly reduce the performance overhead for performing the KNN queries, many scientific applications would benefit from the new method.
+Multimodal retrieval tasks, especially those involving heterogeneous modalities such as text, image (Zhou and Fan 2015), and audio, rely on diverse data sources where conventional indexing techniques such as DBSCAN or $k$-means become ineffective.
+Many real-world multimodal systems, such as those in material science (Rangel DaCosta et al.
+2021), often consist of multiple images and associated textual descriptions, stored in databases that manage structured (tables) and unstructured (blobs) data separately.
+Traditional indexing approaches fail to capture the unified semantic similarity between different modalities, requiring an alternative representation that aligns heterogeneous data within a common space.
+Multimodal machine learning (Li et al.
+2022) provides a means to embed heterogeneous data into a shared vector space (Jia et al.
+2021), enabling unified similarity search.
+However, embeddings from different modalities, such as text and images, are often concatenated, leading to an even higher-dimensional joint representation, which further exacerbates the curse of dimensionality (Aremu, Hyland-Wood, and McAree 2020).
+Given that embeddings generated by models such as BERT (Devlin et al.
+2018) and ViT (Dosovitskiy et al.
+2020) already contain hundreds to thousands of dimensions, concatenation leads to prohibitively high-dimensional representations.
+This dimensionality explosion results in increased storage requirements, computational inefficiency, and degraded nearest-neighbor search performance, particularly in large-scale scientific applications (Wu et al.
+2020; Tshitoyan et al.
+2019).
+Dimensionality reduction offers a potential solution by projecting high-dimensional embeddings into a lower-dimensional space while preserving key relationships.
+Techniques such as PCA (Hotelling 1933) and ISOMAP (Tenenbaum, de Silva, and Langford 2000) aim to minimize the distortion in inter-point distances, facilitating efficient retrieval in reduced spaces.
+However, a fundamental challenge in applying dimensionality reduction to multimodal retrieval is the lack of a principled method to determine the optimal target dimensionality.
+Without an a prior understanding of how well nearest-neighbor structures are preserved post-reduction, practitioners must rely on heuristic or empirical tuning, limiting the robustness of these methods in real-world applications (Zhao et al.
+2023).
+This uncertainty is particularly problematic in tasks such as cross-modal retrieval (Liang et al.
+2024), Retrieval-Augmented Generation (RAG) (Lewis et al.
+2020), and scientific knowledge discovery (Gupta et al.
+2022; Lu et al.
+2023), where the integrity of neighborhood relationships directly impacts retrieval effectiveness.
+These challenges underscore the necessity of understanding the trade-off between dimensionality reduction and nearest-neighbor preservation.
+A framework that enables quantitative estimation of neighbor preservation after dimensionality reduction would allow for more effective multimodal retrieval by mitigating the curse of dimensionality while maintaining computational efficiency.
 Proposed Work
-In order to reduce the dimension of high-dimensional embedding vectors in a meaningful sense, this paper first proposes a new notion, namely Order-Preserving Measure (OPM), for the set of nearest neighbors that is preserved during a dimension-reduction map between two metric spaces.
-In essence, the OPM is formalized as a mathematical function that is well defined over the power-set $\sigma$-algebra over the lower-dimensional metric space.
-This allows us to assign a numerical (real and non-negative) value to an arbitrary subspace of the low-dimensional space and consequently, apply aggregation to quantify the overall accuracy of the dimension-reduction map in terms of preserving the set of $k$-nearest neighbors of each embedding vector.
-Built upon the aforementioned OPM, this paper then constructs a global metric to quantify the closeness of two metric spaces $X$ and $Y$ in terms of, again, the preservation of the set of $k$-nearest neighbors from $X$ to $Y$.
-The objective of constructing such a global metric is to derive a closed-form function among important environmental parameters, such as the dimensionality of metric spaces, the number of data points, and the expected closeness, or accuracy, between the metric spaces in terms of $k$-nearest neighbors.
-The conjectured closed-form function is constructed through multiple observations and qualitative analysis.
-The expected dimension of the lower-dimensional space is positively correlated to both the space’s dimensionality and cardinality (i.e., the number of embedding vectors), but with different weights.
-Specifically, we assign significantly higher weight (exponentiation) on the closeness metric (also known as accuracy, which will be formally defined) than that on the cardinality.
-In addition to theoretical results about the closed-form function, this paper also demonstrates how to incorporate the order-preserving measure and the closed-form function into the workflow of real-world scientific applications.
-We will detail the implementation of both the OPM and the closed-form function, and demonstrate that how they work together with other subsystems, such as dimension-reduction methods (PCA, MDS), neural network models (CLIP, ViT, Bert), and distance metrics (Euclidean, cosine, Manhattan).
-We then evaluate the augmented dimension-reduction system with four multimodal scientific datasets of representative characteristics; our results show that the function accurately captures the intrinsic relationship among the key factors of a dimension-reduction map and therefore, contributes new insights to leveraging advanced AI techniques for complex scientific applications.
+To effectively reduce the dimensionality of high-dimensional embedding vectors while preserving the neighborhood structure, we introduce a mathematically rigorous measure function that quantifies the preservation of nearest neighbors across different metric spaces.
+This measure, formally defined over the power-set $\sigma$-algebra of the lower-dimensional space, quantifies the degree to which $k$-nearest neighbors are preserved.
+By enabling a well-defined aggregation over the entire space, this measure provides a concrete way to evaluate the fidelity of dimension-reduction transformations.
+Building on this measure, we analyze the quantitative relationship between neighborhood preservation and target dimensionality.
+We construct a closed-form function that captures the intrinsic dependencies among key parameters, including the original and reduced dimensionality, the number of data points, and the expected neighborhood preservation accuracy.
+Our analysis reveals that the required lower-dimensional space is positively correlated with both the original dimensionality and dataset cardinality, but their influence varies.
+In particular, neighborhood preservation accuracy exerts an exponentially stronger effect than dataset size, indicating its dominant role in determining the optimal reduced dimensionality.
+In addition to theoretical results, we demonstrate how the proposed measure and closed-form function integrate into practical vision applications.
+We detail their incorporation into dimensionality reduction pipelines using PCA (Hotelling 1933) and MDS (Kruskal and Wish 1978), evaluate their compatibility with state-of-the-art embedding models such as CLIP (Radford et al.
+2021), ViT (Dosovitskiy et al.
+2020), and BERT (Devlin et al.
+2018), and examine their performance under different distance metrics, including Euclidean, cosine, and Manhattan.
+Extensive experiments on multimodal scientific datasets and image-text benchmarks validate the accuracy of the closed-form function in capturing the relationship between dimensionality reduction and neighborhood preservation.
+The results highlight the effectiveness of this framework in mitigating the curse of dimensionality while ensuring robust retrieval performance in vision-based applications.
 Contributions
-In summary, this work makes the following scientific and technical contributions.
-  • We present a new notion for dimension reduction, namely Order-Preserving Dimension Reduction (OPDR), which leads to a more flexible metric that suffices to capture the invariant of $k$-nearest neighbors.
-  • We construct a closed-form function that captures the intrinsic relationship among the target dimensionality and other factors in the context of OPDR.
-  • We implement the proposed OPDR method and demonstrate the effectiveness of OPDR by carrying out an extensive evaluation on real-world multimodal scientific data.
-Preliminaries and Related Work
-Multimodal Scientific Data
-Multimodal scientific data has drawn a lot of research interests.
-For instance, the MELINDA dataset (Wu et al.
+This work makes the following key contributions:
+  • We introduce a mathematically rigorous measure function to quantify the preservation of $k$-nearest neighbors across different dimensional spaces, providing a formal foundation for evaluating dimensionality reduction methods.
+  • We derive a closed-form function that quantitatively describes the relationship between neighborhood preservation and target dimensionality, offering a principled approach to determining the optimal reduced dimensionality.
+  • We empirically validate the proposed framework across multiple datasets and dimensionality reduction techniques, demonstrating that the closed-form function accurately predicts neighborhood preservation and effectively mitigates the curse of dimensionality.
+Background and Related Work
+Multimodal Data Analytics
+Multimodal data analytics has attracted growing research interest across scientific domains.
+For example, the MELINDA dataset (Wu et al.
 2020) focuses on classifying biomedical experiment methods by utilizing multiple data types.
 Similarly, a study on explainable AI (Jin, Li, and Hamarneh 2022) evaluates algorithms within the context of multimodal medical imaging tasks, emphasizing the importance of decision-making based on raw scientific data.
-Researchers have also trained new large language models specifically designed for scientific data.
+In parallel, researchers have developed new large language models specifically designed for scientific data.
 For example, GIT-Mol (Liu et al.
@@
 While these models, along with others such as the unsupervised word embeddings for materials science literature (Tshitoyan et al.
-2019), demonstrate the potential of specialized models in capturing domain-specific knowledge, they are often limited in scope, typically focusing solely on text-based data.
-Multi-modal data is often highly dimensional, making it crucial to find a trade-off between maintain essential information and reducing the data’s complexity.
-Therefore, some scientists have proposed effective methods to address this challenge.
+2019), demonstrate strong potential in capturing domain-specific knowledge, they are often limited in scope, typically focusing solely on text-based data.
+Multimodal data is often highly dimensional, making it crucial to find a trade-off between maintaining essential information and reducing the data’s complexity.
+Given the high dimensionality of multimodal data, researchers have proposed methods to balance information preservation and complexity reduction.
 For instance, the study on Multiscale Feature Extraction and Fusion of Image and Text in VQA (Lu et al.
 2023) introduces advanced techniques to fuse multimodal data, allowing for a more integrated analysis of different data types.
-The research on Semi-Supervised Multi-Modal Learning with Balanced Spectral Decomposition (Hu et al.
+The research on Semi-Supervised Multimodal Learning with Balanced Spectral Decomposition (Hu et al.
 2020) focuses on optimizing information preservation through spectral decomposition, which enhances the ability to capture correlations within the data.
-Additionally, Unsupervised word embeddings in materials science literature (Tshitoyan et al.
+Additionally, Unsupervised word embeddings in materials science (Tshitoyan et al.
 2019) demonstrates the transformation of textual information into vectors, providing a compact and efficient representation of large-scale text data.
@@
 2021), which integrates both text and image data, marking a significant step forward in multimodal data processing.
-All kinds of studies that improve CLIP model has been propsed.
+Various improvements to the CLIP model have been proposed.
 For instance, TiMix (Jiang et al.
@@
 2024) further enhance the alignment between text and images.
-In (Zhang et al.
-2021), authors proposed to convert images into binary codes (hash codes) that preserve image similarity, making it easier to compare images within large-scale datasets.
+In (Zhang et al. 2021), authors proposed to convert images into binary codes (hash codes) that preserve image similarity, making it easier to compare images within large-scale datasets.
 Some prior works proposed methods to perform dimension reduction over the embedding vectors.
@@
 However, the above works assumed that the users of LLMs were able to modify the model structure, which is unrealistic in many scientific applications.
+Multimodal Embedding and Retrieval across Heterogeneous Data
+Recent progress in multimodal representation learning has significantly advanced cross-modal retrieval tasks, which aim to identify semantically aligned items across heterogeneous modalities such as text, images, and audio.
+Traditional vector retrieval methods, including FAISS(Douze et al.
+2024), ScaNN(Guo et al.
+2020), and HNSW (Malkov and Yashunin 2018), enable efficient similarity search for high-dimensional embeddings.
+However, as embedding dimensionality increases, these methods face scalability challenges, leading to degraded retrieval efficiency and increased computational cost.
+While existing approaches focus on scaling similarity search or enhancing retrieval precision, few address the structural integrity of nearest-neighbor relations after dimensionality reduction.
+To address this limitation, our work introduces Order-Preserving Dimension Reduction (OPDR), which reduces embedding dimensionality while preserving $k$-nearest neighbor (kNN) structure, thereby improving retrieval efficiency in large-scale multimodal datasets.
 Order-Preserving Dimension Reduction
@@
 Then, a closed-form function is constructed to quantify the relationship among the space dimensionality, the number of data points, and the OPM.
-Finally, a detailed implementation is presented about how to incorporate the new measure and function into scientific applications.
+Finally, a detailed implementation is presented about how to incorporate the new measure and function real-world multimodal retrieval systems.
 Order-Preserving Measure
-Intuitively speaking, an order-preserving measure (OPM) provides a metric of the number of nearest neighbors that do not change between two spaces.
-For example, if the $2$-closest points (we assume that the spaces are metric spaces such that pair-wise distances are well defined) of $every$ point in a metric space $(X,\delta_{X})$ are still the same $2$-closest points (of each point) in a metric space $(Y,\delta_{Y})$, then we say the map $f:X\to Y$ is order-preserving of 2, or $OP_{2}$.
+Informally, an order-preserving measure (OPM) provides a metric of the number of nearest neighbors that do not change between two spaces.
+For example, if the $2$-closest points (we assume that the spaces are metric spaces such that pair-wise distances are well defined) of everyevery point in a metric space $(X,\delta_{X})$ are still the same $2$-closest points (of each point) in a metric space $(Y,\delta_{Y})$, then we say the map $f:X\to Y$ is order-preserving of 2, or $OP_{2}$.
 We will provide a more formal definition of $OP_{z}$, $z\in\mathbb{Z}^{+}\cup\{0\}$ later; before that, we need to point out a common misunderstanding of this notion regarding inclusiveness, as follows.
 […]
-Mathematically speaking, a measure is a function, say $\mu$, which maps a subset $E_{i}$ of set $X$ in a well-defined $\sigma$-algebra to a value in the extended real line $[0,\infty]$, such that (i) $\mu(\emptyset)$ = 0 and (ii) $\mu\left(\bigcup_{i=1}^{n}E_{i}\right)=\sum_{i+1}^{n}\mu(E_{i})$.
+Formally, a measure is a function, say $\mu$, which maps a subset $E_{i}$ of set $X$ in a well-defined $\sigma$-algebra to a value in the extended real line $0,\infty]$, such that (i) $\mu(\emptyset)$ = 0 and (ii) $\mu\left(\bigcup_{i=1}^{n}E_{i}\right)=\sum_{i+1}^{n}\mu(E_{i})$.
 Because we assume there exists a well-defined $\sigma$-algebra over (the set of) embedding vectors, say $X$, $E_{i}$’s will be understood as disjoint subsets of $X$ unless otherwise stated.
 In fact, we will explicitly construct the $\sigma$-algebra on the target space, $Y$.
-As a side note, a measure function $\mu$ can be also thought of as a homomorphism between the two monoids (i.e., groups without inverses) $(\mathcal{P}(X),\cup)$ and (+$(\mathbb{R}^{+}\cup\{0\},+)$ with kernel $Ker_{\mu}=\{\emptyset\}\subseteq\mathcal{P}(X)$, where $\mathcal{P}(X)$ denotes the power set of $X$.
+As a side note, a measure function $\mu$ can be also thought of as a homomorphism between the two monoids (i.e., groups without inverses) $(\mathcal{P}(X),\cup)$ and $(\mathbb{R}^{+}\cup\{0\},+)$ with kernel $Ker_{\mu}=\{\emptyset\}\subseteq\mathcal{P}(X)$, where $\mathcal{P}(X)$ denotes the power set of $X$.
 However, our analysis in the following will not require any group-theoretical results.
@@
 Let $Y$ denote the target (i.e., low-dimensional) metric space (we omit the metric functions here) of the dimension-reduction map.
-In practice, the cardinality of $Y$ is finite (i.e., there is an upper limit of number of embedding vectors in scientific applications), which is of course countable.
+In practice, the cardinality of $Y$ is finite (i.e., there is an upper limit of number of embedding vectors in vision applications), which is of course countable.
 The $\sigma$-algebra $\mathcal{M}_{Y}$ of $Y$ is simply the power set of all the mapped vectors in $Y$, i.e., $\mathcal{M}_{Y}=\mathcal{P}(Y)$, which is obviously a $\sigma$-algebra on $Y$ because $Y$ is countable and $\mathcal{P}(Y)$ is closed for any countable number of set unions.
-After constructing its $\sigma$-algebra, we are ready to define the measure on $\mathcal{M}_{Y}$.
-Let $E^{Y}_{k,i}$ denote the set of the $k$-nearest neighbors of $y_{i}$, where $y_{i}\in Y$.
-The same notation can be defined for space $X$, which denotes the original metric space where the original vectors of $y_{i}$ reside, i.e., $x_{i}=f^{-1}(y_{i})$, where $f$ denotes the dimension-reduction function.
-Let $\mu$ be a function defined as follows,
+After constructing its $\sigma$-algebra, we are ready to define the measure on $\mathcal{M}_{Y}$. Let
+$E^{Y}_{k,i}$ denote the set of the $k$-nearest neighbors of $y_{i}$, where $y_{i}\in Y$.
+The same notation can be defined for space $X$, which denotes the original metric space where the original vectors of $y_{i}$ reside,
+i.e., $x_{i}=f^{-1}(y_{i})$, where $f$ denotes the dimension-reduction function. Let $\mu$ be a function defined as follows,
 $\mu(F_{j})=\frac{\left|F_{j}\cap E^{Y}_{k,i}\cap E^{X}_{k,i}\right|}{k},$ (1)
-where $F_{j}\in\mathcal{M}_{Y}$. It should be clear that $\mu\in[0,1]\subset[0,+\infty]$ because $\left|E^{Y}_{k,i}\right|\in[0,k]$, where ||$\cdot$|| ⋅ | denotes the cardinality of a set.
+where $F_{j}\in\mathcal{M}_{Y}$. It should be clear that $\mu\in[0,1]\subset[0,+\infty]$ because $\left|E^{Y}_{k,i}\right|\in[0,k]$, where $|\cdot|$ denotes the cardinality of a set.
 We argue that the function defined in Eq.
@@
 To prove property (ii), it suffices to show that if $F_{j}=F_{1}\cup F_{2}$ and $F_{1}\cap F_{2}=\emptyset$, then $\mu(F_{1}\cup F_{2})=\mu(F_{1})+\mu(F_{2})$; this is because $\mathcal{M}_{Y}$ is finite and the above binary relationship can be extended finite times.
-To simply the notation, let $E=E^{Y}_{k,i}\cap E^{X}_{k,i}$.
+To simplify the notation, let $E=E^{Y}_{k,i}\cap E^{X}_{k,i}$.
 It follows that $\mu(F_{j})$ = $\frac{\left|(F_{1}\cup F_{2})\cap E\right|}{k}$ = $\frac{\left|(F_{1}\cap E)\cup(F_{2}\cap E)\right|}{k}$.
@@
 The reasoning is that the $k$-nearest neighbors can be thought of the discrete representation of the neighborhood of the referred point, or vector.
-If we extend the idea to a continuous counterpart, i.e., given a point $p$ and its neighborhood or an open set around it, say $\epsilon(p)$, the question becomes under what conditions the open set of the mapped vectors, namely $f(p)$, corresponds to the open set of $p$ in a dimension-reduction function $f$.
+If we extend the idea to a continuous counterpart, i.e., given a point $p$ and its neighborhood or an open set around it, say $\epsilon(p)$, the question becomes under what conditions the open set of the mapped vectors,
+namely $f(p)$, corresponds to the open set of $p$ in a dimension-reduction function $f$.
 In topology and analysis, such a function carrying the open sets in the forward direction is called an open map, which is unfortunately not the same (in fact, the inverse) condition under which the function is continuous.
@@
 The previous section defines an additive measure $\mu$ on the $\sigma$-algebra of the projected space $Y$; This section investigates the relationship between $\mu$ and other parameters, such as the space dimensionality and the number of data points, or space cardinality.
-The idea is that we will construct a function for the above variables that we conjecture are critical during the dimension reduction, largely based on our experiment observations and mathematical intuition.
-We will later in the evaluation section verify the constructed function as our hypothesis.
+The idea is to construct a function for the above variables which we hypothesize to be critical for dimensionality reduction, largely based on our empirical observations and mathematical intuition.
+We will later verify the constructed function as a working hypothesis in the evaluation section.
 We assume that we have an existing dimension-reduction method on hand and we are interested in quantifying the target dimensionality of the lower-dimensional space in which the set of $k$-nearest neighbors is an invariant, as defined in Eq.
@@
 $\text{A}^{X}_{k}(Y)=\frac{1}{m}\cdot\sum_{i=1}^{m}\frac{\mu_{i}\left(Y\setminus\{y_{i}\}\right)}{k},$ (2)
-where $m=\left|Y\right|=\left|X\right|$, $y_{j}\in Y$, and $u_{i}(\cdot)$ is the measure on $Y$ over $y_{i}$ as defined in Eq.
+where $m=\left|Y\right|=\left|X\right|$, $y_{j}\in Y$, and $\mu_{i}(\cdot)$ is the measure on $Y$ over $y_{i}$ as defined in Eq.
 (1).
@@
 The parameters in our hypothetical function include the accuracy $A_{k}$ between two metric spaces $X$ and $Y$, the number of data points $m=\left|X\right|=\left|Y\right|$, and the dimensionality of the spaces.
-In practice, the dimension of the domain space $X$ is determined by the neural network model; therefore, our function would assume that $X$’s dimension is a constant and only involve the dimension of $Y$, denoted by $dim(Y)$.
+In practice, the dimension of the domain space $X$ is determined by the neural network model; therefore, our function would assume that $X$’s dimension is a constant and only involves the dimension of $Y$, denoted by $dim(Y)$.
 Thus, we expect to construct a function $g$ as follows,
@@
 It follows that the real-world users can simply compose the functions $g$ and $f$, i.e., $f\circ g$, to ensure that the set of $k$-nearest neighbors is an invariant between two spaces $X$ and $Y$.
-The construction of $g$ is as follows. We conjecture the form of $g$ based on the following observations.
+The construction of $g$ is based on the following observations:
   • Firstly, $dim(Y)$ is positively influenced by both the accuracy $A_{k}$ and the cardinality $m$.
@@
 On the other hand, a small change of the number of data points may or may not significantly change the distribution of the space.
-Based on the above discussion, we conjecture that the function $g$ takes the following form
+Based on the above discussion, we postulate that the function $g$ takes the following form
 $g(a,b)=\mathcal{O}(b\cdot 2^{a}),$ (3)
@@
 (4).
-Incorporation into Scientific Data Management
-We will first describe how the embedding vectors are generated from multimodal scientific data and then present how we implement the hypothetical function and incorporate it into popular dimension-reduction methods.
-To convert scientific data into embedding vectors using CLIP (Contrastive Language-Image Pretraining), Vision Transformer (ViT), and BERT (Bidirectional Encoder Representations from Transformers), which are all transformer-based models, we begin by collecting and preparing the scientific data, which includes both textual data from HDF5 files and images in formats such as TIFF, PNG, or JPEG.
-The textual data is extracted by reading the Hierarchical Data Format (HDF5) files, where the dataset names and contents are concatenated into a single string representation.
-For instance, the HDF5 file used in our study contains a dataset with dimensions of (139, 139, 92) along with supporting metadata, providing a structured representation of the scientific data.
-For the image data, each image file is read and converted to RGB format.
-The images, including results obtained at four different resolutions: HAADF, BF, ADF, and ABF, are found to have consistent dimensions of (369, 369) pixels.
-Additionally, a summary image with dimensions of (941, 905) pixels was included.
-Then, we generate embeddings for both textual and image data using the CLIP, ViT, and BERT models.
-For textual data, we utlized BERT model to generate embbedings, which processes the concatenated textual information.
-Image embeddings are produced by the ViT model, which processes the image files.
-Then, we utilized CLIP model to analyze both textual and image data.
-The generated embeddings maintain their default dimensionality, which is specific to each model: BERT produces embeddings of 768 dimensions, while ViT also outputs 768-dimensional embeddings.
-For CLIP, the textual data is embedded via CLIP’s text encoder, resulting in 512-dimensional embeddings.
-Similarly, the image data is processed through CLIP’s image encoder, also yielding 512-dimensional embeddings.
-To create a comprehensive representation of the multimodal scientific data, we combined these embeddings by concatenating the 512-dimensional text and image vectors into a single, unified embedding vector of 1024 dimensions.
-All generated embeddings, derived from both textual data and image data at various resolutions, are collected and stored for further analysis.
+Integration into Multimodal Data Retrieval
+We will first describe how the embedding vectors are generated from multimodal data, specifically the image-text pair and then present how we implement the hypothetical function and incorporate it into popular dimension-reduction methods.
+To leverage transformer-based models, including CLIP (Contrastive Language-Image Pretraining), Vision Transformer (ViT), and BERT (Bidirectional Encoder Representations from Transformers), for converting multimodal data into embedding vectors, we conducted experiments on three distinct datasets: materials science data, Flickr30k, and OmniCorpus.
+Each dataset consists of multimodal data, incorporating text from HDF5 files, natural language text, and images in TIFF, PNG, or JPEG formats.
+The experimental results demonstrate that our approach maintains strong applicability across different domains, including both scientific data and natural images.
+The generated embeddings retain the default dimensionality of each respective model: BERT and ViT both produce 768-dimensional embeddings.
+For CLIP, the text and image encoders each output 512-dimensional vectors.
+We construct unified multimodal representations by directly concatenating embeddings from different modalities—for example, combining CLIP text and image vectors into a single 1024-dimensional vector.
+In the case of audio-text data from the ESC-50 dataset, we generate embeddings using BERT (768D) for textual labels and PANNs CNN14 (Kong et al.
+2020) (2048D) for audio signals.
+These are concatenated into 2816-dimensional joint vectors.
+All embeddings—spanning image, text, and audio modalities—are stored for subsequent dimensionality reduction and retrieval analysis.
 Following the embedding vector extraction, we focus on the dimensionality reduction phase.
 We evaluated several mainstream techniques including PCA, and MDS, among which PCA consistently outperformed others in terms of maintaining the integrity of location information in our datasets.
-Our investigations revealed a notable correlation between the effectiveness of PCA in preserving the spatial relationships and the ratio $\frac{n}{m}$, where $n=dim(Y)$ and $m=\left|Y\right|$, of the target dimension to the number of samples used in the PCA process.
+Our investigations revealed a notable correlation between the effectiveness of PCA in preserving the spatial relationships and the ratio
+$\frac{n}{m}$, where $n=dim(Y)$ and $m=\left|Y\right|$, of the target dimension to the number of samples used in the PCA process.
 We adopted various regression models to elucidate the relationship between the accuracy of preserved relative location information and the ratio $\frac{n}{m}$.
@@
 Evaluation
-This section evaluates the effectiveness of the proposed method for KNN-preserved dimension reduction on multiple scientific data sets.
-To ensure the robustness of our method across various scenarios, we evaluate the method with three Transformer-based models for extracting embedding vectors, two popular dimension-reduction techniques, and three distance metrics.
+This section evaluates the effectiveness of the proposed method for KNN-preserved dimension reduction on multiple data sets.
+To ensure the robustness of our method across various scenarios, we evaluate the method with several Transformer-based models for extracting embedding vectors, two popular dimension-reduction techniques, and three distance metrics.
 All results suggest that the proposed method is highly effective.
@@
 Data Sets
-The dataset used for evaluation were sourced from the Material Project  (The Materials Project Accessed 2024) and categorized into four distinct datasets: experimental Observable, stable, metal, magnetic.
-To be more specific, experimental observable dataset includes 33,990 data points, stable dataset consists of 48,884 data points, metal dataset has 72,252 data points, and magnetic dataset consists of 81,723 data points.
+The evaluation was conducted using seven datasets: four scientific datasets from the Materials Project (The Materials Project Accessed 2024), two natural image-text pair datasets (Flickr30k (Young et al.
+2014) and OmniCorpus-037 CC (Li et al.
+2024)), and one audio-text dataset (ESC-50 (Piczak 2015)).
+Specifically, the observable, stable, metal, and magnetic datasets contain 33,990; 48,884; 72,252; and 81,723 data points, respectively.
+Flickr30k includes 31,014 image-text pairs, while OmniCorpus-037 CC provides 3,878,063 such pairs.
+The ESC-50 dataset comprises 2,000 environmental audio clips, each paired with a textual label describing the sound event (e.g., “dog barking”, “siren”, “rain”).
 Platform
-For our experiments, we chose CloudLab (Duplyakin et al.
-2019) as the testing platform.
-In this experiment, the specific machine specifications we selected are as follows: All of our experiments are carried out on a CloudLab machine with the following specifications: Node Type: c6420.
-The machine is equipped with two sixteen-core Intel Xeon Gold 6142 CPUs at 2.6 GHz, 384 GB ECC DDR4-2666 Memory, and two Seagate 1TB 7200 RPM 6G SATA hard drives.
-The network interface card (NIC) is a dual-port Intel X710 10GbE.
-We have installed the following libraries in conda environment: transformer, sklearn, torch, pandas, numpy, os, pickle, h5py, shutill, matlotlib, and pyprismatic.
+Our experiments were carried out on the NSF-sponsored computing platform CloudLab (Duplyakin et al.
+2019), specifically the d7525 nodes with the following specifications.
+The machine is equipped with two 16-core AMD EPYC 7302 CPUs at 3.00GHz, 128GB ECC Memory (8×16 GB 3200MT/s RDIMMs), and two 480 GB 6G SATA SSDs along with one 1.6 TB PCIe4 x4 NVMe SSD for storage.
+The network interface card (NIC) is a dual-port Mellanox ConnectX-6 DX 100Gb NIC, with one port supporting 200Gb connectivity for experimental use.
+Additionally, the machine is equipped with an NVIDIA 24GB Ampere A30 GPU.
+We have installed the following libraries in conda environment: transformer, sklearn, torch, pandas, numpy, os, pickle, h5py, shutil, matplotlib, and pyprismatic, PIL.
 OPDR on Various Data Sets
-We first extracted embedding vectors from four datasets with distinct characteristics (Observable, Unstable, Metal, Non-magnetic) using the CLIP model.
+We first use the CLIP model to extract embedding vectors from six datasets, comprising four scientifically curated multimodal datasets (Observable, Unstable, Metal, Non-magnetic), two widely adopted vision-language benchmarks (Flickr30k and OmniCorpus), as well as the ESC-50 dataset (Piczak 2015) for audio-text representation.
 Subsequently, we computed the distances between vectors using the L2-norm and performed dimensionality reduction using PCA.
 To simplify notation, we let $n=dim(Y)$, denote the dimension of the lower-dimensional space.
-For these four groups of datasets, we further divided them into eight subsets each, with sample sizes $m$ where $m\in\{10,20,30,40,50,60,70,80\}$, to test whether the distribution patterns of data points were consistent across different sample sizes.
+For these four groups of material datasets, we further divided them into eight subsets each, with sample sizes $m$ where $m\in\{10,20,30,40,50,60,70,80\}$.
+For the remaining two multimodal datasets, we divided them into five subsets each, with sample size $m$ where $m\in\{10,50,100,150,300\}$, to test whether the distribution patterns of data points remain consistent across different sample sizes.
 Observations from Figure 6 to Figure 6 indicate a strong positive correlation between accuracy and the ratio of $n$ to $m$.
@@
 Figure 4: Non-magnetic
-Through comparative analysis of the four data sets discussed above, we observe that our proposed OPDR algorithm performs consistently across multiple data sets and exhibits the anticipated data patterns, validating our initial hypotheses.
+Figure 5: Flicker30k
+Figure 6: OmniCorpus
+Through comparative analysis of the six data sets discussed above, we observe that our proposed OPDR algorithm performs consistently across multiple data sets and exhibits the anticipated data patterns, validating our initial hypotheses.
 In the remainder of this section, we will focus on the Observable Material dataset, exploring the effects of using different Transformer-based models, dimensionality reduction techniques, and various distance metrics.
 Influence of Embedding Models
-We continued our experiments on the ‘Observable’ dataset, merely changing the model used for extracting embedding vectors.
-We employed three distinct transformer-based models: BERT (Bidirectional Encoder Representations from Transformers), Vision Transformer (ViT), and CLIP (Contrastive Language-Image Pre-training) for processing and analyzing scientific data.
-Specifically, the BERT model was used to analyze text data extracted from HDF5 files, ViT was utilized for processing image data in TIFF and PNG formats, and CLIP was applied to analyze comprehensive multimodal datasets.
-This enabled us to achieve more holistic and unified representations of scientific data.
-Through multiple experiments with these three models, we obtained data fit lines for each model to represent the relationship between accuracy and the ratio $n/m$.
-All three fit lines indicated that as $n$ approaches $m$, our accuracy first increases rapidly and then converges slowly.
-The results from the three different models were consistent with our initial hypotheses.
-Furthermore, we observed that the obtained fit lines were essentially overlapping, indicating that the choice of model does not alter the data structure.
-Utilizing these diverse transformer-based models, we were still able to ensure the accuracy and consistency of data relationships.
-Using different models does not change the data format; thus, in practical applications, it is essential to consider the properties of the data and its intended use to select the most appropriate model for experimentation.
-Figure 5: Data Fitting for Different Embedding Models
-Figure 5 shows that using different models does not alter much the accuracy.
+To evaluate the impact of different embedding models, we conducted further experiments on the Observable dataset, alongside Flickr30k and OmniCorpus from the multimodal datasets, using different models for extracting embedding vectors.
+We employed three distinct transformer-based models—BERT (Bidirectional Encoder Representations from Transformers), Vision Transformer (ViT), and CLIP (Contrastive Language-Image Pre-training)—to process and analyze the datasets.
+This enabled us to obtain more holistic and unified representations across different data domains.
+Through extensive experiments with these three models, we derived data fit lines to represent the relationship between accuracy and the ratio $n/m$.
+All three fit lines consistently indicated that as $n$ approaches $m$, accuracy initially increases rapidly and then gradually converges.
+The results from the three different models aligned with our initial hypotheses.
+Additionally, for datasets with structured relationships—such as material-related datasets—the obtained fit lines were nearly overlapping, suggesting that the choice of model does not significantly alter the underlying data structure.
+In contrast, for the Flickr30k and OmniCorpus datasets, differences between models were more pronounced, though the overall trend still followed the expected logarithmic pattern.
+This indicates that while different transformer-based models may introduce variations in representation, they maintain the fundamental structure of the data.
+These findings reinforce that, regardless of the chosen model, the core data relationships remain stable.
+Thus, in practical applications, selecting the most appropriate model should be guided by the characteristics of the dataset and the specific task requirements, rather than concerns over fundamental data integrity.
+Figure 7: Data Fitting for Embedding Models on Material Dataset
+Figure 8: Data Fitting for Embedding Models on Flicker dataset
+Figure 9: Data Fitting for Embedding Models on OmniCorpus dataset
+Figure 10: Data Fitting for Dimension Reduction on Material Dataset
+Figure 11: Data Fitting for Dimension Reduction on flicker dataset
+Figure 12: Data Fitting for Dimension Reduction on OmniCorpus dataset
+Figure 12 to Figure 12 show that using different models does not influence much the accuracy.
 This implies that the closed-form function is applicable to a variety of neural network models.
 Therefore, in practical applications, one should consider the nature of the data and its intended use to select the most appropriate model for experimentation.
-The original data from which the regression line is calculated can be found in Figure 8 in the Appendix.
 Influence of Dimension-Reduction Methods
-The scientific data analyzed in this study exhibits extremely high dimensionality, often reaching up to hundreds of thousands of dimensions.
-Despite the application of advanced embedding techniques, the resulting vectors still contain several hundred to several thousand dimensions, posing considerable challenges for effective data analysis and interpretation.
-To tackle these issues, we employed two principal dimensionality reduction methods—Multidimensional Scaling (MDS) and Principal Component Analysis (PCA)—both renowned for their ability to efficiently reduce data complexity while maintaining the integrity of the relationships among data points.
-Given that the majority of contemporary dimensionality reduction techniques build on the foundational theories of PCA and MDS, our study specifically focused on these methods to evaluate their practical efficacy.
-By implementing MDS and PCA, we aimed to simplify the dataset in a way that preserves the relative positioning of data points, thereby facilitating more manageable and interpretable analysis.
-By comparing the performance of PCA and MDS on the same dataset in Figure 6, we observe that PCA is more sensitive to changes in $n/m$ and can reach higher accuracy more quickly.
-Moreover, PCA can achieve a maximum accuracy of 100%, surpassing the peak accuracy attainable by MDS.
-This demonstrates PCA’s superior performance in preserving the relative positions of data points.
-Additionally, employing different dimensionality reduction techniques does not alter the overall data pattern; the relationship between accuracy and $n/m$ still aligns with our hypothesis.
-Figure 6: Data Fitting for Different Dimension-Reduction Methods
-From Figure 6, although the overall data pattern remains constant and both PCA and MDS align with our prior hypotheses, the choice of dimensionality reduction technique is found to have a noticeable effect on the fitting outcomes.
+The datasets analyzed in this study exhibit extremely high dimensionality, often reaching hundreds of thousands of dimensions.
+Even after applying advanced embedding techniques, the resulting vectors typically retain several hundred to several thousand dimensions, presenting significant challenges for effective analysis and interpretation.
+To address these issues, we employed two principal dimensionality reduction methods—Multidimensional Scaling (MDS) and Principal Component Analysis (PCA)—both widely recognized for their ability to efficiently reduce data complexity while preserving the structural integrity of data relationships.
+As many modern dimensionality reduction techniques are fundamentally derived from PCA and MDS, our study specifically focused on these methods to assess their practical effectiveness.
+By comparing the performance of PCA and MDS on the same dataset in Figure 12, we found that PCA is more sensitive to changes in $n/m$ and converges to higher accuracy more quickly.
+Additionally, in the materials science dataset, PCA achieves a peak neighborhood preservation accuracy of 100%, surpassing the highest accuracy attainable by MDS.
+In the other two multimodal datasets, PCA also demonstrates greater precision compared to MDS, further validating its superiority.
+Moreover, while different dimensionality reduction techniques may affect specific numerical values, they do not alter the overall data pattern; the relationship between accuracy and $n/m$ remains consistent with our hypothesis.
+From Figure 12 to Figure 12, although the overall data pattern remains constant and both PCA and MDS align with our prior hypotheses, the choice of dimensionality reduction technique noticeably affects the fitting outcome.
 In practical applications, it is crucial to weigh the specific contexts and experimental data at hand.
 This assessment will inform the selection of the most appropriate dimensionality reduction method to achieve the desired level of accuracy.
-The original data from which the regression line is calculated can be found in figure 9 in the Appendix.
-Influence of Distance Metrics
-Figure 7: Data Fitting for Different Distance Metrics
-In this section, we explore the impact of using different distance metrics—namely L1 norm, L2 norm, and cosine distance—on the Observable dataset and CLIP model.
-The geometric meanings of these metrics are distinct:
-  • L1 norm (Manhattan distance): Measures the sum of the absolute differences of their coordinates.
-It represents a very intuitive distance metric, similar to the shortest path between two points in a grid-like structure.
-  • L2 norm (Euclidean distance): Calculates the straight-line distance between two points.
-It is the most commonly used distance metric, applicable in a wide range of natural and engineering problems.
-  • Cosine distance: Measures the cosine of the angle between two vectors, reflecting the difference in direction rather than magnitude, often used to gauge similarity in text or high-dimensional data.
-Experimental observations shown in Figure 7 revealed that as $n/m$ increases, the behaviors of L2 norm and cosine distance are similar, whereas L1 norm performs slightly worse than the other two.
-Despite the change in distance metrics not altering the overall data pattern—they still show the trend proposed in Eq (4)—the choice of distance metric remains crucial for specific scenarios.In practical use, experimenters should consider the application scenarios of their data to decide which distance calculation method to use.
-The original data from which the regression line is calculated can be found in figure 10 in the Appendix.
 Conclusion and Future Work
-This paper solves an open problem in multimodal scientific data management regarding the high dimensionality of embedding vectors generated by large language models.
-Theoretical results include new notions (both point-wise and space-level) to measure the closeness of two metric spaces in terms of their $k$-nearest neighbors and a closed-function to reveal the relationship between space dimensionality, space cardinality, and space similarity (or, accuracy of the dimension-reduction map).
-The closed-form function is implemented and incorporated into other subsystems, such as dimension-reduction methods (PCA, MDS), neural network models (CLIP, ViT, Bert), and distance metrics (Euclidean, cosine, Manhattan), which collectively enable a new dimension-reduction approach for efficiently manage and analyze multimodal scientific data.
-Our future work will dive deeper into the theoretical foundation of the closed-form function among metric spaces’ dimensionality, cardinality, and accuracy of the dimension-reduction map; the fact that the proposed order-preserving measure is a well-defined measure may shed more insight than we originally thought.
-We also plan to investigate the extensibility of the proposed method for production vector database systems, such as PostgreSQL and pgvector.
-Appendix A Additional Experimental Results
-According to Figure 8, the data distributions obtained from the three Transformer-based models are highly similar.
-In contrast, the distribution from the BERT model is more dispersed, while the distributions from CLIP and ViT models are more concentrated.
-The results indicate that the CLIP model performs the most stably, with the overall patterns of the three models being consistent, aligning with our initial hypotheses.
-The fitting lines for the three groups of data points have been discussed previously in Figure 5.
-Figure 8: Original Data for Neural Network Models
-Figure 9 clearly shows the accuracy distributions for the ratio $n/m$ using two different dimensionality reduction techniques, PCA and MDS.
-Both PCA and MDS exhibit highly concentrated data points.
-Comparatively, PCA demonstrates higher accuracy as $n$ approaches $m$, indicating that PCA has stronger performance capabilities.
-Overall, the data patterns of both methods are similar, aligning with our initial hypotheses.
-The fitting lines for these two dimensionality reduction methods have already been discussed in figure 6.
-Figure 9: Original Data for PCA and MDS
-Figure 10 clearly illustrates the accuracy distributions for the ratio $n/m$ using different distance metrics.
-Data points generated by the L2-norm and cosine distance highly overlap, indicating similar performance of these metrics.
-In contrast, the L1-norm shows lower sensitivity and fails to reach the precision of other methods.
-Overall, the three datasets exhibit similar patterns: accuracy rapidly increases as $n$ approaches $m$ and then slowly converges to a peak.
-This observation aligns with our initial hypothesis.
-The fitting lines for the three sets of data points obtained using different distance metrics have been discussed in figure 7.
-Figure 10: Original Data for Distance Metrics
+This paper proposes a new method, namely Order-Preserving Dimension Reduction (OPDR), to address the challenge of high-dimensional embeddings in multimodal data analytics, particularly those generated by large language models.
+From a theoretical perspective, this work introduces new concepts (both point-wise and space-level) to measure the closeness of two metric spaces in terms of their $k$-nearest neighbors and a closed-form function to characterize the relationship among dimensionality, cardinality, and similarity (i.e., the accuracy of the dimension-reduction map).
+The closed-form function is incorporated into a practical ecosystem by complementing other system components, such as dimension-reduction methods (PCA, MDS) and models (CLIP, ViT, Bert).
+Extensive evaluation demonstrates both the effectiveness and practical utility of OPDR.
+Our future work will further explore the theoretical foundation of the closed-form function, particularly the relationship among metric spaces’ dimensionality, cardinality, and the accuracy of the dimension-reduction map.
+The fact that the proposed order-preserving measure is a well-defined metric may provide deeper insights.
+Additionally, we plan to investigate the extensibility of our method in production vector database systems, such as PostgreSQL and pgvector, to support large-scale retrieval tasks.
+Beyond multimodal retrieval, we aim to explore potential applications in medical imaging, where high-dimensional embeddings from MRI (Chamberland et al.
+2019; Sridhar et al.
+2022), CT scans, and vital records could benefit from dimension reduction while preserving similarity relationships—enabling efficient, accurate, and scalable content-based retrieval for medical diagnosis.